the difference between the length and the width of a rectangle is 2 units. the perimeter is 40 units. write and solve a system of equations to determine the length and width of the rectangle.
Answer by NancyLam(35)

Very easy
perimeter rectangle is :
(L+W)x2=40
L+W= 40/2
L+W =20
L= w+2 (difference between is 2)
2+w + w =20 (subtitution L= 2+W)
2w+2=20
w= (20-2)/2
w=9
Check:W=9,L=11
(9+11)x2= (20)(2)=40
I hope it help